OpenStudy (anonymous):

find the zeros of the polynomial function f(x)=x^2-4x-60

OpenStudy (anonymous):

answers -10, 6 -14,5 -6, 10 -5, 14

OpenStudy (rulnick):

0 = x^2 - 4x - 60 = (x-10) (x+6) So the zeros are at the values where x-10=0 and x+6=0. Can you solve from there?

OpenStudy (anonymous):

so... -6,10?

OpenStudy (rulnick):

Yes!

OpenStudy (anonymous):

ok tank you! :)

OpenStudy (rulnick):

welcome